OpenStudy (unicornpoopcookies):

Find the slope & y intercept of 5x - 6 y =0

OpenStudy (unicornpoopcookies):

@Austin1617

Directrix (directrix):

5x - 6 y =0 must be written in the y = m*x + b form where m is the slope and b is the y-intercept.

Directrix (directrix):

I'll move on given that you want to work with Austin1617. No worries.

OpenStudy (unicornpoopcookies):

I know im supposed to subtract the 5x add it to the other side and then divide by 6 right

OpenStudy (anonymous):

@UnicornPoopCookies

OpenStudy (anonymous):

Your slope is 5/6 and y-intercept is 0

OpenStudy (unicornpoopcookies):

Thank you!

OpenStudy (anonymous):

Next questions?

Directrix (directrix):

>Your slope is 5/6 and y-intercept is 0 How did you get that? Please explain. Thanks. @Austin1617

OpenStudy (anonymous):

@Directrix You should look at my math question.

Directrix (directrix):

What about the answer you posted in this thread. How did you get it?

Directrix (directrix):

I see a graph but no explanation for the OP.

OpenStudy (unicornpoopcookies):

is it positive or negative

OpenStudy (unicornpoopcookies):

positive right

OpenStudy (anonymous):

-6y=-5x Divide each term in -6y=-5x by -6 \[-\frac{ 6y }{ -6}=-\frac{ 5x}{ -6 }\] Factor 6 out of -6 \[\frac{ 6y }{ 6\times-1 }=-\frac{ 5x }{ -6 }\] Next you will cancel out the common factor. 6 is the common factor. \[\frac{ y }{ 6\times-1 }=-\frac{ 5x }{ -6}\] Rewrite the expression. \[\frac{ y }{ -1 }=-\frac{ 5x}{ -6 }\] Move the negative one form the denominator of y/-1. \[-1 \times y=-\frac{ 5x }{ -6}\] \[-(-1\times y)=-\frac{ 5x }{ -6 }\] \[-(-y)=-\frac{ 5x }{ -6 }\] \[-1\times-1y=-\frac{ 5x }{ -6 }\] \[1y=-\frac{ 5x }{ -6 }\] \[y=-\frac{ 5x }{ -6 }\]

OpenStudy (anonymous):

\[y=-(-\frac{ 5x }{ 6 })\] \[y=-1\times -1(\frac{ 5x }{ 6 })\] \[y=1(\frac{ 5x }{ 6 })\] \[y=\frac{ 5x }{ 6 }\] Slope-intercept formula y=mx+b m=5/6 b=0

OpenStudy (anonymous):

@Directrix @UnicornPoopCookies